OpenStudy (anonymous):

Solve for x: 4x - 16y = 32

OpenStudy (kendricklamar2014):

What have you done so far @ninamae05

OpenStudy (anonymous):

4x - 16y = 32 +16 y = + 16 y ---------------------- 4x =16 y + 32 ------------------ 4 x= 4y + 8

OpenStudy (anonymous):

I am not sure @KendrickLamar2014

OpenStudy (kendricklamar2014):

That is correct :)